If a man ran 1/5 of mile in 12 minutes ,what is his speed in miles per hour?

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

If a man ran 1/5 of mile in 12 minutes ,what is his speed in miles per hour?
-----
[(1/5)/12](miles/minute)(60 min/1 hr)
----
= (1/5)*1*60)/(12*1*1)(miles/hr.
----
= 12/12 miles hr
----
= 1 mph
======================
cheers,
Stan H.

Answer by Alan3354(69443)

If a man ran 1/5 of mile in 12 minutes ,what is his speed in miles per hour?
--------
12 mins = 12/60 hour = 1/5 hour
----
d = r*t
r = d/t = (1/5)/(1/5)
r = 1 mi/hr
============
I wouldn't call that running.
Slow crawling, maybe.